# A Generative Approach to LLM Harmfulness Detection with Special Red Flag Tokens

### v3

@@ Abstract
Most safety training methods for large language models (LLMs) are based on fine-tuning that forces models to shift from an unsafe answer to refusal when faced with harmful requests. Unfortunately, these drastic distribution shifts generally compromise model capabilities. To avoid that, we propose to expand the model’s vocabulary with a special token we call red flag token () and propose to train the model to insert this token into its response at any time when harmful content is generated or about to be generated. Our approach offers several advantages: it enables the model to explicitly learn the concept of harmfulness while marginally affecting the generated distribution, thus maintaining the model’s utility. It also evaluates each generated answer and provides robustness as good as adversarial training without the need to run attacks during training. Moreover, by encapsulating our safety tuning in a LoRA module, we provide additional defenses against fine-tuning API attacks.
@@ 1 Introduction
To make large language models (LLMs) that are robust against a determined adversary, practitioners rely on several security layers such as model hardening via fine-tuning (Zou et al., 2024; Xhonneux et al., 2024; Sheshadri et al., 2024), perplexity filters (Alon & Kamfonas, 2023), or harmfulness classifiers (Inan et al., 2023; Sharma et al., 2025). However, as the model capabilities progress, so does their attack surface, as innate abilities can be used to circumvent defences (Huang et al., 2024)—e.g., using a low-resource language to jailbreak the model. It is therefore natural to embed safety mechanisms into the models themselves, which will scale with these capabilities. Ideally, an LLM would have no harmful faculties, but this is unrealistic as many desirable skills can be both useful and damaging depending on the context. Therefore, we argue that the model itself should detect when it is being used harmfully and that this behaviour can be complementary to standard safety training whose goal is to remove harmful faculties from the model.
To address this, we propose a method to detect harmful behaviour using the LLM’s generative process, which does not compromise its helpfulness. Our approach uses an additional special token for LLMs to predict when the model considers that its capabilities are being used unsafely. We call this a red flag token () and train the model to output this token at any time during the generation of a harmful response, while not changing its response after or in benign contexts. This special token is excluded from the user vocabulary and can be filtered from streamed responses. This complementary layer of safety has several benefits. First, our method only requires the to be included in the answer with a single token change, while other methods such as adversarial training like Xhonneux et al. (2024), force models to completely change their output distribution from a harmful response to a refusal. Secondly, our method does not rely on a specific safe answer, meaning that if the model is broken, e.g., through pre-filling (Andriushchenko et al., 2024) or random sampling (Huang et al., 2023), we can still output a to tag the answer as harmful and have it be filtered.
We demonstrate the feasibility of this approach by designing a loss function consisting of three components: (i) a cross-entropy loss on generating the (ii) a Kullback-Leibler (KL) loss term on the generation after the (see Figure 1) and (iii) a KL loss term on benign utility conversations. We test our approach with several adversarial attacks on three open-source models: Llama3.2-3B-IT (Grattafiori, 2024), Mistralv3-IT (Jiang et al., 2023), and Phi-3.5 (Haider et al., 2024).
Finally, inspired by the idea of task arithmetic (Ilharco et al., 2023), we introduce the concept of storing the safety aspect of the model in a LoRA module (Hu et al., 2021) and show that it can be applied to retain detection capabilities after an attacker has leveraged a fine-tuning API to remove the safety training of the model.
To sum up, the contributions of this paper are:
@@
we propose a specialised red flag token to reliably detect harmfulness at each generation step, even under strong adversarial attacks, including pre-filling, sampling, and automated jailbreaks like GCG and PAIR;
we show empirically that our approach generalises beyond the training distribution, effectively handling significantly longer inputs than those seen during training;
Finally, we demonstrate that our safety module can be efficiently stored in a LoRA module to be applied to detect harmful outputs from harmfully fine-tuned models.
2 Method
2.1 Threat model
We assume that LLM access is gated behind some web interface or API with no access to model weights, logits, or direct control of input/output processing—commonly referred to as a black box setting. The main assumption we make is that the service provider can evaluate the logits and output of the model before streaming it to the user, including filtering tokens such as assistant tokens or our token. We also consider a more permissive gray box setting where the user may have access to extra features such as pre-filling, and also some white box settings where the user has access to model weights. Finally, we consider the most permissive setting, a fine-tuning attack setting where the user has access to some fine-tuning API.
2.2 Our Loss
We assume that we have a dataset of harmful prompt-continuation pairs and a data set of harmless (a.k.a., benign) pairs. To train the model to flag unsafe generations, we sample an index from a distribution defined over the positions of the harmful continuation , where is a minimum offset that avoids flagging too early in the generation. The continuation is then split into three parts: the prefix , the red flag token , and the suffix . We use to denote the cross entropy and to denote the Kullback-Leibler divergence (KL). We consider our reference model . Our loss consists of three components: first, to ensure our model outputs the red flag token in harmful completions, we use a standard language modeling cross-entropy loss on all harmful completion tokens starting at the minimum offset up to and including the token:
To maintain model performance and reduce distribution shift as much as possible without increasing the likelihood of a harmful answer, we use a KL divergence on the tokens after the :
All these losses put together, we get:
Note that none of the loss functions make a harmful continuation more likely, enabling our approach to be complementary to other safety fine-tuning techniques. Figure 1 summarises our approach, with the training procedure fully defined in Algorithm 1.
3 Experiments
3.1 Models & Datasets
We fine-tune Llama3.2-3B-IT (Grattafiori, 2024), Mistralv3-IT (Jiang et al., 2023), and Phi-3.5 (Haider et al., 2024) using our algorithm on the Harmbench (Mazeika et al., 2024) training set using 32 adversarial continuations per harmful prompt generated via the ablation attack (Arditi et al., 2024) along with original refusals. We use 5000 samples from the Alpaca dataset (Taori et al., 2023) as benign prompts. All models listed have a set of reserved special tokens as part of their tokenizers, allowing us to avoid extending the vocabulary and instead we re-purpose one of these unused special tokens to be the .
We assess model utility on standard LLM benchmarks, MMLU (Hendrycks et al., 2021), ARC-E and ARC-C (Chollet, 2019), as well as a Harmless dataset of benign prompts consisting of 119 random prompts from Ultrachat200k (validation split), and 40 benign prompts with a similar syntax as Harmbench provided by (Xhonneux et al., 2024, Appendix I). We also use this benign dataset to compute the false positive rate of refusal (or ) in Figure 3.
For adversarial robustness evaluation, we compute the defence success rate (DSR) of different attacks on the Harmbench Standard test set (Mazeika et al., 2024) that contain 159 harmful prompts, balancing with an equal number of benign prompts. Either a refusal or surpassing a certain probability threshold counts as a successful defence. The thresholds are set per model and are , , and for Llama3.2-3B-IT, Mistralv3-IT, and Phi-3.5, respectively.
All models are trained on a single A100 GPU with LoRA (Hu et al., 2021) and a batch size of 64 using the AdamW optimizer (Loshchilov & Hutter, 2017) (for more hyper-parameters see the Appendix A).
3.2 Design decisions
There are several design decisions to consider:
The cross-entropy loss on can be computed on each index before and including the sampled position or only on . In other words, we have the choice to allow the model for flexibility of when to output at the cost of potentially overfitting more because we now train the model to output immediately after the instruction token. In particular, this forces the model to judge the prompt quite strongly, leading to a higher probability for in a refusal as well. In practice, we tested both approaches and saw better results computing the cross entropy up to and including index . A potential solution to avoid over-fitting after the instruction token is to have a minimum offset into the harmful continuation both for sampling as well as the cross-entropy term, which we apply to Mistralv3-IT.
The sampling distribution of the index at which to insert the is a key choice. We tested both a geometric distribution as well as a uniform distribution over the harmful continuation. We decide to use a geometric distribution, which means that fewer positions will be used in the cross-entropy loss, and more will be used in .
The attention mask in the can also be amputated not to include the harmful prompt and the harmful continuation before the such as to force the model to store the harmful continuation information in the embedding. In other words, with probability , We choose to apply this trick probabilistically with probability 0.5, such as to make training more stable but still encourage the to be meaningful for generation.
Calibrating sensitivity can be done in several ways. We can add a logit bias term to the before passing the logits through a softmax, modifying the distribution to be:
where is a logit for the token and is a user-specified bias to re-weight . We can also directly monitor without adding a logit bias term or looking for explicit generations of , and set a model-dependant threshold in which to flag the generation as potentially harmful. We elect to use the second strategy as it works well and is simpler to apply in practice.
3.3 Baselines
We consider three baselines. Since the models come with safety training, we record the natural refusal rate. We then consider the refusal rate of CAT (Xhonneux et al., 2024)—an adversarial training technique using continuous attacks, where we replicate their training procedure using our datasets for a fair comparison. Finally, we adapt Jain et al. (2024) to our setting, whereby we insert the at the first position of the assistant’s response and call this baseline fixed-position-rf.
3.4 Robustness Evaluation
We compute the defence success rates of the following attacks:
Sampling Attacks jailbreak models by sampling the response multiple times (Hughes et al., 2024). Occasionally, the model may eventually provide an answer to a harmful prompt after repeated queries. In our experiments, we sample up to 16 times or until the model responds, as evaluated by the official classifier for text behaviours in HarmBench000huggingface.co/cais/HarmBench-Llama-2-13b-cls. We use a temperature of and a top- value of for the sampling attack.
Pre-filling and sampling are gray-box attacks using additional features such as temperature-based sampling and manually injecting text into a model generation. GCG is a white-box attack, and PAIR is a black-box attack. An attack is successful if the model does not refuse or the does not surpass a model-specific threshold. We evaluate refusals using GPT-4o (OpenAI et al., 2024) as a judge.
Figure 2 summarises our results. We first observe that our approach maintains near-perfect utility across all models. In particular, all baselines achieve utility scores comparable to the base model, with CAT showing the lowest score on Harmless, and this allows for a good comparison point in terms of robustness utility trade-off. CAT results in a 4% drop on the Harmless evaluation that indicates over-refusal.
Our proposed approach is able to nearly perfectly defend against the gray-box attacks, i.e., Sampling and Pre-filling across all models. We find that on Llama3.2-3B-IT and with the Pre-filling attack, the approach results in a significantly higher DSR than CAT, which mainly fails to refuse most harmful queries when a pre-filled affirmative response is present. The fixed-position-rf is marginally worse than on all attacks, with a larger gap on GCG.
We also report very good DSR against PAIR, and some improvement against GCG (with respect to the base model) for the approach. In general, comparing the base model and , we find that even when requests are not being refused, we can still defend against the malicious attack with the detection of .
3.5 Fine-tuning attacks
We consider the fine-tuning attack threat model discussed in Section 2.1, where the attacker is able to fine-tune the model by submitting training data through an API. We train a LoRA module that stores the weight changes for the model to insert the in harmful continuations on Llama3.2-3B-IT. We can then apply this LoRA module one or more times after the attacker has fine-tuned the base model and see if is still generated during harmful generations. For the fine-tuning attack we also use a LoRA module and use the data and hyper-parameters from Qi et al. (2024), which consists of about 100 harmful examples and continuations.
As baselines we consider both the aforementioned CAT (Xhonneux et al., 2024) as well as the fixed-position-rf stored in LoRA modules. In addition, we test whether applying each of these approaches multiple times can further improve robustness. Finally, we check whether we can combine the LoRA modules for the approach with the orthogonal CAT approach.
As before, we use the same Harmbench test set and the same Harmless dataset from Figure 2; the results are in Figure 3. This figure demonstrates the fundamental robustness and utility trade-off that exists in terms of false positive rate (FPR) and true positive rate (TPR)—i.e., defence success rate. The goal is to improve the Pareto front. The LoRA maintains good performance in harmful contexts, with the fixed-position baseline performing poorly. The adversarial training baseline (CAT) improves safety but cannot be calibrated, producing no Pareto front. However, applying the CAT LoRA once or twice yields alternative robustness trade-offs. Unlike CAT, our LoRA module allows fine-grained calibration and a strong Pareto front. We also show that combining the CAT and modules yields complementary benefits, improving robustness and enabling calibrated trade-offs between utility and safety (TPR/FPR).
We also validate that on benign fine-tuning our module does not impact the gained performance significantly. We test this on GSM8k (Cobbe et al., 2021) in chat mode under strict match of the answer—more details is provided in Appendix C.
3.6 Generalisation to Longer Contexts
During training, we only include data with a single turn of user/assistant interaction, with being injected following some distribution biased towards the start of the first assistant’s response. We validate that our approach generalises to multiple user/assistant interactions without over-fitting to the start of the conversation by sampling a number of benign conversations from UltraChat-200k, and then appending harmful queries from our test set. By monitoring the probability of throughout the model’s generation, we find that the prediction of does not deteriorate with increased sequence length. One chat interaction is shown in Figure 4 with Figure 6 containing the corresponding text; the probabilities consistently remain low in the regime in which it was trained (during the assistant’s turn) and increases sharply in the presence of harmful content during pre-filling and generation. Aggregate statistics are shown in Figure 5.
3.7 Limitations
Expectedly, our approach does not defend very well against attacks that have complete access to the model weights such as GCG, although it still achieves a considerable improvement in robustness compared to the base model. This shows that the association between harmful generation and the is circumvented by these attacks, demonstrating a limitation of our approach which we plan to investigate further. Furthermore, in the fine-tuning attack setting, there is still a trade-off between robustness and utility in the face of an attacker, albeit a much stronger Pareto-front than without the module.
Another limitation of our method is the amount of hyperparameters that are introduced that require tuning. While it is not too difficult to achieve a decent trade-off, one can likely achieve much better performance with better hyperparameter tuning as we did not have the resources to effectively tune the model (all the trainings were done with a single GPU). In addition, our ability to associate the with harmfulness relies on the data provided.
@@ 4 Related Work
Beyond standard pre-training, LLMs are typically trained with preference optimization techniques such as RLHF (Ouyang et al., 2022) or DPO (Rafailov et al., 2023) to be more aligned with human preferences. Jailbreaks can be incorporated into this preference alignment phase to increase resilience to such attacks (as is often done with red-teaming methods), but this does not often generalise to novel jailbreaks. Historically, in the context of vision models, actively training against adversarial attacks in an online manner (i.e., adversarial training) is the only method that has shown increased adversarial robustness (Madry et al., 2017). However, in the context of language, most discrete attacks are prohibitively expensive to use online. Mazeika et al. (2024) train against adversarial suffixes generated by GCG, but continually update a pool of examples rather than generate each attack from scratch. Other approaches perform adversarial training by attacking the embedding or latent space of the model (Xhonneux et al., 2024; Sheshadri et al., 2024) which is much more efficient to compute and transfers to discrete attacks. Beyond adversarial training, newer defences target and alter harmful representations in order to prevent a model from producing harmful outputs entirely (Zou et al., 2024). Independent from training a model to be more robust to jailbreaks is to classify and judge the potential harmfulness of the generated text, often with another LLM fine-tuned for this task (Inan et al., 2023; Feuer et al., 2024), although this does require additional resources to classify the outputs. Concurrent work Huang et al. (2025) has shown that classifiers alone are often not sufficient, further making the case that other approaches are needed especially against permissive but common threat models such as the fine-tuning box attack.
@@
Several works have explored training or utilising special tokens for specific purposes. Burtsev et al. (2020) prepend “memory” tokens to an input prompt on a target task. Goyal et al. (2023) append “pause” tokens, which are hypothesised to give the LLM a buffer sequence to reason over before producing an output. Mu et al. (2023) train LLMs to compress longer prompts into smaller sets of “gist” tokens as a means to shorten the context. Xiao et al. (2023) prepend “attention sinks” to improve generalization to long-context sequences. LLMs have also been trained to use a variety of tools (such as a calculator or internet access), which are denoted and invoked via special tokens (Schick et al., 2023).
@@
Most closely related to our approach is the recent work of Jain et al. (2024), where a model is trained to prefix an output with a special refusal or response token based on the behaviour of whether the model refuses or responds to a prompt. While their approach is related in that special tokens are leveraged in the context of alignment, the approach and objective are conceptually different. Their method correlates these tokens with behaviour (i.e., refusal or response) in order to better calibrate such behaviours, whereas our approach correlates a special token with some implicit notion of a concept (i.e., harmfulness), without modifying the model’s original behaviour. This conceptual difference leads to drastically different losses in the formulation. For instance Jain et al. (2024) do not propose a KL divergence with a reference model (Equation 2) to maintain the predictions similar to the reference model after is outputted which hurt the model’s utility and is not complementary with standard safety training (instead it is a way to calibrate the model post-hoc safety training). Moreover, their model is only trained to output a “behavioural token" (e.g., “refuse" or “respond") at the beginning of the answer, which is significantly less efficient to detect harmfulness, as shown in our experiments. In contrast, our work proposes an approach that is complementary to standard safety training where the model essentially acts as an “implicit judge” on its own generated output, improving its transparency and providing a clear signal to evaluate potentially harmful generations without incurring any additional computational cost at inference time.

### v4

# A Generative Approach to LLM Harmfulness Mitigation with Red Flag Tokens

@@ Abstract
Many safety post-training methods for large language models (LLMs) are designed to modify the model’s behaviour from producing unsafe answers to issuing refusals. However, such distribution shifts are often brittle and degrade performance on desirable tasks. To address these pitfalls, we propose augmenting the model’s vocabulary with a special “red flag” () token, and training the model to insert this token whenever harmful content is generated or imminent. This approach enables the model to explicitly learn the concept of harmfulness in its representations, with minimal impact on utility due to the marginal change in the generated distribution of natural language. Moreover, because the token is embedded in the model’s vocabulary, we can naturally leverage the LLMs’ generalization capabilities, such as in-context learning (ICL) and out-of-distribution generalization to languages that are not formally supported (e.g., Japanese for Llama3). In particular, we demonstrate that through ICL alone, the model can learn to initiate reflective reasoning upon generating the token at inference, which steers the response away from harmful continuations or enables self-correction when the flag is raised falsely. This approach is orthogonal and complementary to existing safety technique (such as safety classifiers or standard safety training) and easier to evaluate in comparison to natural language refusals, as it does not require a human or automated judge to assess the harmlessness of the answers.
@@ 1 Introduction
Defending large language models (LLMs) against adversarial users relies on multiple security layers, including safety fine-tuning (Zou et al., 2024; Xhonneux et al., 2024; Sheshadri et al., 2024), perplexity filters (Alon & Kamfonas, 2023), or harmfulness classifiers (Inan et al., 2023; Sharma et al., 2025). However, as model capabilities advance, so does their attack surface, as innate abilities can be used to circumvent defences (Huang et al., 2024)—e.g., using a low-resource language to jailbreak the model. It is therefore natural to embed safety mechanisms directly into models, allowing defences to scale with capabilities. While an LLM with no harmful faculties would be ideal, this is unrealistic: many desirable skills can be either beneficial or harmful depending on context. We therefore argue that models themselves should be able to detect when their capabilities are being misused, complementing standard safety training that aims to remove harmful behaviours.
To address this, we propose a method to detect harmful behaviour through the LLM’s generative process without compromising helpfulness. Our approach uses an additional special “red flag” () token that the model learns to output when detecting unsafe capability usage. We train the model to output this token at any time during the generation of a harmful response, while not changing its response after the token, or in benign contexts. This special token is excluded from the user vocabulary and can be filtered from streamed responses. At inference time, the token serves as a flexible signal of potential harmfulness. As a defence, we consider two settings: filtering the entire response and replace it with a safe alternative (Section˜4.2), or triggering safety-oriented reflective reasoning before providing an answer (Section˜4.3.1) as illustrated in Figure˜1.
This approach offers several technical advantages. First, it requires minimal distribution shift: only a single token insertion rather than forcing complete output rewrites from harmful responses to refusals, as in traditional safety training and adversarial training methods (Xhonneux et al., 2024; Casper et al., 2024). This single explicit token also provides a natural and principled target for adversarial training, as we can directly optimize adversarial perturbations to suppress generation while encouraging affirmative responses, and then train against these attacks to improve robustness. Second, because the method is agnostic to specific safe responses, the model can still flag harmful content even when compromised through pre-filling (Andriushchenko et al., 2024) or random sampling (Huang et al., 2023). This also enables straightforward evaluation, since the detection of a token provides an objective signal without requiring human or automated judges to assess response harmlessness.
Third, unlike external safety classifiers that evaluate complete responses post-hoc, integrating the harmfulness signal directly into generation enables finer-grained detection and allows the model to condition on its own internal safety assessments. This integration with the generative process allows seamless application of existing LLM techniques: we demonstrate in Section˜4.3.1 that through prompting and in-context learning (ICL) alone, models can learn to use tokens as a soft signal for reflective chain-of-thought reasoning—evaluating whether the flagged content is truly harmful or not—rather than as a hard filter, enabling recovery from false positives on benign inputs. Finally, our approach demonstrates good generalization capabilities across the LLM’s natural capabilities; we show that our method generalizes effectively to other languages including supported languages like Spanish, and unsupported languages like Japanese (on our particular models), despite our fine-tuning dataset primarily comprising English prompts and completions (Section 4.3.2).
To summarise, the contributions of this paper are:
@@
We propose to use a special red flag token to detect harmfulness at each generation step, even under strong adversarial attacks, including pre-filling, sampling, and automated jailbreaks like GCG and PAIR;
We demonstrate the feasibility of this approach by designing a loss function consisting of three components: (i) a cross-entropy loss on generating the token (ii) a Kullback-Leibler (KL) loss term on the generation after the token and (iii) a KL loss term on benign utility conversations;
We extend our training algorithm to an adversarial training paradigm to further improve results;
We show that our approach generalizes beyond the training distribution, including (a) integration with natural language instructions via in-context learning, where the model learns to use the token as a soft signal for reflective safety reasoning, and (b) cross-lingual transfer, including unsupported languages.
@@ 2 Related Work
A few recent works have proposed filtering harmful data from the pretraining phase (Li et al., 2025; O’Brien et al., 2025; Chen et al., 2025b). Our approach suggests an alternative to filtering such data, that could be implemented by simply tagging harmful data during pre-training. In addition to some filtering and rephrasing, contemporary work (Maini et al., 2025) explores the idea of “tagging” harmful content with a special token (corresponding to the use of a red flag token during pre-training) and demonstrates how it can be used to filter potential generations during beam search. In contrast, our work demonstrates how such a special red flag token can be learned at post-training time and how it can be used to either detect harmful content or be used to trigger reasoning, leveraging the generalization capabilities of LLMs.
Beyond standard pre-training, LLMs are typically trained with preference optimization techniques such as RLHF (Ouyang et al., 2022) or DPO (Rafailov et al., 2023) to be more aligned with human preferences. Jailbreaks can be incorporated into this preference alignment phase to increase resilience to such attacks (as is often done with red-teaming methods), but this does not often generalize to novel jailbreaks (Andriushchenko et al., 2024). Historically, in the context of vision models, actively training against adversarial attacks in an online manner (i.e., adversarial training) is the only method that has shown increased adversarial robustness (Madry et al., 2017). However, in the context of language, most discrete attacks are prohibitively expensive to use online. Mazeika et al. (2024) train against adversarial suffixes generated by GCG, but continually update a pool of examples rather than generate each attack from scratch. Other approaches perform adversarial training by attacking the embedding or latent space of the model (Xhonneux et al., 2024; Sheshadri et al., 2024) which is much more efficient to compute and transfers to discrete attacks. Beyond adversarial training, newer defences target and alter harmful representations in order to prevent a model from producing harmful outputs entirely (Zou et al., 2024). Independent from training a model to be more robust to jailbreaks is to classify and judge the potential harmfulness of the generated text, often with another LLM fine-tuned for this task (Inan et al., 2023; Feuer et al., 2024; Sharma et al., 2025), although this does require additional resources to classify the outputs. Huang et al. (2025) has shown that classifiers alone are often not sufficient, further making the case that other approaches are needed.
@@
Several works have explored training or utilising special tokens for specific purposes. Burtsev et al. (2020) prepend “memory” tokens to an input prompt on a target task. Goyal et al. (2023) append “pause” tokens, which are hypothesised to give the LLM a buffer sequence to reason over before producing an output. Mu et al. (2023) train LLMs to compress longer prompts into smaller sets of “gist” tokens as a means to shorten the context. Xiao et al. (2023) prepend “attention sinks” to improve generalization to long-context sequences. Chen et al. (2025a) append “defensive tokens” before the LLM input to defend against prompt injection attacks. LLMs have also been trained to use a variety of tools (such as a calculator or internet access), which are denoted and invoked via special tokens (Schick et al., 2023).
@@
Closely related to our approach is the recent work of Jain et al. (2024), where a model is trained to prefix an output with a special refusal or response token based on the behaviour of whether the model refuses or responds to a prompt. While their approach is related in that special tokens are leveraged in the context of alignment, the approach and objective are conceptually different. Their method correlates these tokens with behaviour (i.e., refusal or response) in order to better calibrate such behaviours, whereas our approach correlates a special token with some implicit data-driven notion of a concept (i.e., harmfulness), without modifying the model’s original behaviour. This conceptual difference leads to drastically different losses in the formulation. For instance Jain et al. (2024) do not propose a KL divergence with a reference model (Equation˜2) to maintain the predictions similar to the reference model after a token is generated, since their objective was to provide a way to calibrate the model’s refusal sensitivity rather than to provide an additional safety layer. Moreover, their model is only trained to output a “behavioural token" (e.g., “refuse" or “respond") at the beginning of the answer, which is significantly less efficient to detect harmfulness, as shown in our experiments. In contrast, our work proposes an approach that is complementary to standard safety training where the model essentially acts as an “implicit judge” on its own generated output, improving its transparency and providing a clear signal to evaluate potentially harmful generations without incurring any additional computational cost at inference time.